Use for 6-10: Below are hypothetical data on lung cancer cases in a given year in a population. Consider smoking to be the exposure and cancer to be the outcome of interest. Number of lung Smoking Status Smokers Nonsmokers Total cancer cases Population 40,000 60,000 50 100,000 6. 7. 8. 9. Calculate the relative risk. Calculate the odds ratio. Calculate the risk difference. Assuming the smoking is the cause of the excess lung cancer in smokers, calculate and interpret the attributable fraction among the exposed. 10. Assuming the smoking is the cause of the excess lung cancer in smokers, calculate and interpret the population attributable fraction.
Answer #1

Question 6

Here probability of lung cancer in nonsmokers = 6/60000 = 0.0001

probability of lung cancer in smokers = 44/40000 = 0.0011

relative risk = 0.0011/0.0001 = 11

Question 7

Here odds to catch lung cancer in nonsmokers = 6/(60000 - 6) = 0.0001

odds to catch lung cancer in smokers = 44/(40000 - 44) = 0.0011

Odd Ratio = 0.0011/0.0001 = 11.011

Question 8

Risk difference = 0.0011 - 0.0001 = 0.001

Question 9

Attritutable fraction among the exposed = (0.0011 - 0.0001)/0.0011 = 0.9091 or 90.91%

Question 10

Population attributable fracion = (proportion of cases exposed) x (attributable proportion in the exposed)

= [p * (RR - 1)/{p(RR-1) + 1)]

= [0.4 * (11 - 1)/(0.4 * 10 + 1)]

= 4/5

= 0.8